# Identifying Sparse Treatment Effects

### v1

@@ 4 A Baseline Approach
Lin’s estimator [Lin, 2013] improves the CUPED estimator by running separate regressions in the treated and control groups, with the agnostic property that it does not induce higher variance than the difference-in-mean estimator without any model assumptions. Lin’s estimator estimates $\tau$ as
@@
To adapt the baseline approach to use pre-treatment covariates, instead of the simple difference-in-mean estimator, one can use covariates-adjusted difference-in-mean estimators. For either CUPED or Lin’s method ($\text{method}\in\{\text{CUPED},\text{Lin}\}$), we consider the following objective:
$\hat{\beta}=\operatorname*{argmin}_{\beta}\Bigg{|}\Bigg{|}\left(\frac{{\hat{\tau}_{\text{method}}}_{j}}{{\sqrt{\hat{\Sigma}_{\text{method}}}_{jj}}}\right)_{j=1}^{p}-\beta\text{ }\Bigg{|}\Bigg{|}_{2}^{2}\quad\text{subject to }||\beta||_{0}\leq s_{\tau}.$ (13)
@@
With these definitions, inference proceeds again as in Section 3 using the second split with covariates-adjusted $\tilde{Y}$, but restricted to the set $\hat{S}_{\tau}$.
Under some regularity and model assumptions (as in Section 5.2), this baseline approach would consistently recover $S_{\tau}$ as $n\xrightarrow[]{}\infty$ (similarly as in Section 5.2). In our numerical simulations in Section 6.1, however, we see that the baseline approach may not work well in low signal-to-noise-ratio (SNR) regimes (see Figure 3). Notably, many sensor data examples outlined in Section 1 fall within such low SNR regimes. In the next section, we propose subset selection methods using the Lasso, which is known to generally perform well in low SNR regimes [Hastie et al., 2020].
@@ 5 Our Proposed Approach:
Our approach starts from the following observation: Rather than considering each individual outcome representation separately, they can be considered as a group as represented by the Hotelling’s $T^{2}$ test statistic in (5). This leads us to consider the following objective:
@@
where $\hat{\tau}^{S}_{\text{DiM}}$, $\hat{\Sigma}_{\text{DiM}}^{S}$ on the right-hand side are computed for each set $S$.
In the next subsection, we show that in fact the preceding problem is equivalent to fitting a weighted linear regression with carefully chosen weights $W_{i}$ as
$\hat{\beta}=\operatorname*{argmin}_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta})^{2}\quad\text{subject to }||\beta||_{0}\leq s_{\tau},$ (15)
where $||\beta||_{0}=\sum_{i=1}^{p}1\{\beta_{i}\neq 0\}$ is the $\ell_{0}$ norm of $\beta$ and $\hat{S}_{\tau}=\{j|\hat{\beta}_{j}\neq 0\}$. Note that we have treatment assignments as response variables and centered outcomes as explanatory variables. Here, $\bar{Y}$ is the mean of outcomes on the pooled data.
Now that we have reformulated the problem as a weighted linear regression with subset selection, we can employ various statistical tools such as forward subset selection, backward subset selection, and Lasso. In the following, we focus on Lasso.
Unlike the objective (7) which has a closed-form solution, solving (15) often presents computational complexity challenges. To avoid this issue, we develop our method using Lasso. The Lasso solves a convex relaxation of (15) where we replace the $\ell_{0}$ norm with the $\ell_{1}$ norm as
$\hat{\beta}=\operatorname*{argmin}_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta})^{2}\quad\text{subject to }||\beta||_{1}\leq t$
where $||\beta||_{1}=\sum_{i=1}^{p}|\beta_{i}|$, and $t\geq 0$ is a tuning parameter. Due to convex duality, the above problem is equivalent to the more common penalized form
$\hat{\beta}=\operatorname*{argmin}_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta})^{2}+2\lambda||\beta||_{1}$ (16)
@@
Having identified a subset $\hat{S}_{\tau}$ on the first data split, inference can proceed as in Section 3 using the second data split, but restricted to the set $\hat{S}_{\tau}$. We can specialize the definitions in (1) and (4) to the subset $\hat{S}_{\tau}$, and carry out inference using the second data split as suggested in (3) or (6) defined with these $\hat{S}_{\tau}$-specific quantities. Note that in (3), we would still apply a correction factor for multiple testing, but this will be significantly smaller after our subset selection step when $|\hat{S}_{\tau}|\ll p$.
@@
$(\hat{\beta},\hat{\alpha})=\operatorname*{argmin}_{\beta,\alpha}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta}-(X_{i}-\overline{X})^{\intercal}{\alpha})^{2}+2\lambda||\beta||_{1}$ (17)
where the $\ell_{1}$ penalty is on $\beta$ and $\lambda\geq 0$ is a tuning parameter. Then, $\hat{S}_{\tau}=\{j|\hat{\beta}_{j}\neq 0\}.$ Note that
$\operatorname*{argmin}_{\beta}\min_{\alpha}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta}-(X_{i}-\overline{X})^{\intercal}{\alpha})^{2}$
$=\operatorname*{argmin}_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(\tilde{Y}-\bar{\tilde{Y}})^{\intercal}\beta)^{2},$
@@
Having identified $\hat{S}_{\tau}$, using the second split, we compute the covariate-adjusted difference-in-means estimator such as the CUPED estimator or Lin’s estimator on $\hat{S}_{\tau}$. The inference proceeds again as in Section 3 with covariates-adjusted $\tilde{Y}$ restricted to the set $\hat{S}_{\tau}$.
Our method shares a close relationship with the projection test [Lopes et al., 2011, Huang, 2015, Liu et al., 2022] in high-dimensional mean testing, as both involve finding sparse $\beta$ using a quadratic objective function with $\ell_{1}$ penalization. We focus on causal inference, and in particular as a result, we also consider pre-treatment covariates. Moreover, we employ a linear regression specification, making it easier for practitioners to implement our method with tools like glmnet in R [Friedman et al., 2010].
Algorithm 1 provides a summary of the proposed method.
Input: $(T_{i},X_{i},Y_{i})_{i=1}^{n}$
@@
1 Randomly split the data set into two folds $\mathcal{D}_{1}=\{(T_{i},X_{i},Y_{i})\}_{i=1}^{n_{1}}$ and $\mathcal{D}_{2}=\{(T_{i},X_{i},Y_{i})\}_{i=1}^{n_{2}}$
@@ 2 Using the first split $\mathcal{D}_{1}$, estimate
where $\tilde{Y}_{j}$ is defined as in (11) or (12), and estimate its variance $\hat{\Sigma}$ using (4) with $\tilde{Y}$ on $\hat{S}_{\tau}$.
Algorithm 1 Single-Split Algorithm
Extension to Multi-Sample Splitting.
Methods relying on sample splitting are known to be sensitive to the arbitrary choice of a one-time random split of the data [Meinshausen et al., 2008]. Therefore, it is recommended to employ multiple splits and aggregate p-values from each split, as outlined in Meinshausen et al. [2008]. This improves the robustness of the method by reducing the dependence on a single split. In Algorithm 2, we present an algorithm box where we apply the multi-sample splitting results from Meinshausen et al. [2008] to our proposed method. Note that one consequence of using multiple splits is that while we obtain an aggregated p-value for testing the null hypothesis, we no longer obtain an estimated subset localizing the treatment effect (since each split produces one such subset). Aggregating the subsets remains an interesting direction for future work.
Input: $(T_{i},X_{i},Y_{i})_{i=1}^{n}$, $\gamma$
Output: valid p-values under $H_{0}:\tau=0$
1 for b = 1, …, B do
3       Calculate p-value:
  1. (a)
    (P-values with multiple testing) The p-value for each $Y_{j}$ is
    $p_{j}^{(b)}$ $=1\quad\text{for }j\notin\hat{S}_{\tau}^{(b)},$
    where $\Phi$ is the cdf of the standard Gaussian distribution.
  2. (b)
    (P-value with Hotelling $T^{2}$ statistic) The p-value is
Aggregate p-values as
Algorithm 2 Multi-Split Algorithm
In the remainder of the section we first discuss the reduction of (14) to a weighted least squares regression; and then present a result that provides conditions under which we recover $S_{\tau}$.
@@ 5.1 Reduction to Weighted Linear Regression
Consider the following weighted least squares regression problem:
$\hat{\beta}=\arg\min_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta})^{2}.$ (19)
We set the weights so that the least-squares problem becomes equivalent to a signal strength maximization problem, where the signal strength is defined using Hotelling $T^{2}$ statistic in (5).
We define $Y^{S}$ to be an outcome vector restricted to the set $S$ as $Y^{S}:=(Y_{j}|j\in S)^{\intercal}.$
@@
On the right-hand side, ${\hat{\tau}_{\text{DiM}}^{S}}$ and $\hat{\Sigma}_{\text{DiM}}^{S}$ are obtained for each $S$. Note that the right-hand side is a Hotelling $T^{2}$ statistic in (5). This means that with our chosen weights, the weighted least-squares problem becomes equivalent to maximization of the Hotelling $T^{2}$ statistic, as desired.
@@ 5.2 Recovery Results for the Lasso
Recall that $\hat{\beta}$ is defined as
$(\hat{\beta},\hat{\alpha})=\operatorname*{argmin}_{\beta\in\mathbb{R}^{p},\alpha\in\mathbb{R}^{m}}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}\beta-(X_{i}-\overline{X})^{\intercal}\alpha)^{2}+2\lambda_{n}||\beta||_{1}$ (26)
@@
Define the (population-level) covariate-adjusted potential outcomes $Z$ as
$Z:=Y-(1-\pi)\cdot\left[\text{Var}(X)^{-1}\text{Cov}(X,Y(1))\right]^{\intercal}X-\pi\cdot\left[\text{Var}(X)^{-1}\text{Cov}(X,Y(0))\right]^{\intercal}X.$
Note that this adjustment of outcomes $Y$ on pre-treatment covariates $X$ resembles pre-treatment covariates adjustments with linear regressions in Lin [2013] and Jin and Ba [2023]. If pre-treatment covariates are not available, set $Z=Y$.
We will provide the error bound for $||\hat{\beta}-\beta^{*}||_{1}$, where the optimal solution of the population risk $\beta^{*}$ is defined as
@@
By Sherman-Morrison formula,
@@
In our setting in Section 3, the treatment effect $\tau$ is sparse, with only $s_{\tau}$ nonzero elements. In the following, we assume that $\beta^{*}$ is sparse, with only $s^{*}$ nonzero elements. This is assuming that the inverse of $\Sigma_{Z}$ is sparse. Below, we provide examples where the inverse of ${\Sigma_{Z}}$ and thus $\beta^{*}$ are sparse under an appropriate structural causal model [Bollen, 1989, Pearl, 2009]. Let $S^{*}=\{j|\beta^{*}\neq 0\}$. In examples below, we identify setting where $S_{\tau}\subseteq S^{*}$. Note that for the remainder of this section, we present the recovery results for $S^{*}$, which would include $S_{\tau}$ in settings below. We also provide simulation results when the sparsity assumption is slightly violated in the Appendix C.
@@
Assume that $\epsilon_{1},\dots,\epsilon_{p}$ are mean zero and uncorrelated with each other. Then, we have
$S_{\tau}:=\{j:\tau_{j}\neq 0\}=S^{*}:=\{j:\beta^{*}_{j}\neq 0\}.$
@@
In this example we consider the same linear structural causal model as in the previous example, but instead of assuming that $\epsilon_{1},\dots,\epsilon_{p}$ are uncorrelated with each other, we assume a cluster structure: there exist clusters $C_{1},\dots,C_{c}$ such that if $j\in C_{i}$ and $j^{\prime}\in C_{i^{\prime}}$ with $i\neq i^{\prime}$, then $\epsilon_{j}$ and $\epsilon_{j^{\prime}}$ are uncorrelated. Then, when $S_{\tau}\subseteq C_{1}$, we have
$S_{\tau}\subseteq S^{*}=C_{1}.$
With a sparse cluster structure, $s^{*}=|C_{1}|\ll p$.
@@
Continue to assume the same linear structural causal model, but now assume that $\epsilon$ follows a multivariate normal distribution and “most” variables are conditionally independent of each other given the remaining variables. For example, we could consider an autoregressive model (AR($k$)) for time-series data. Then, the inverse covariance matrix of $\epsilon$ is sparse, and we have
$S_{\tau}\subseteq S^{*}=\text{Markov Blanket of }S_{\tau}\text{ under the graph of }\epsilon.$
With a sparsely connected graph structure of $\epsilon$, $s^{*}\ll p$.
For our theoretical results, we assume sub-Gaussian designs for $X$, $Y(0)$, $Y(1)$, and thus for $Z(0),Z(1)$ as well. Then, we add assumptions about their moments and the invertibility of the covariance matrix of $X$.
@@
$X_{1},\dots,X_{n}$ are independent and identically distributed sub-Gaussian vectors with mean $\mu_{X}$ and covariance $\Sigma_{X}$. Moreover, $Y_{1}(0),\dots,Y_{n}(0)$ and $Y_{1}(1),\dots,Y_{n}(1)$ are independent and identically distributed sub-Gaussian vectors with mean $\mu^{0}_{Y}$ and $\mu^{1}_{Y}$ respectively and covariance $\Sigma^{0}_{Y}$ and $\Sigma^{1}_{Y}$ respectively. Additionally, $Z_{1}(0),\dots,Z_{n}(0)$ and $Z_{1}(1),\dots,Z_{n}(1)$ are independent and identically distributed sub-Gaussian vectors. Let the mean be $\mu^{0}_{Z}$ and $\mu^{1}_{Z}$ respectively and the covariance be $\Sigma^{0}_{Z}$ and $\Sigma^{1}_{Z}$ respectively. More specifically, each variable $(Z(0))_{j}/\sqrt{(\Sigma^{0}_{Z})_{jj}}$ is sub-Gaussian with parameter $\sigma_{Z}^{0}$ and $(Z(1))_{j}/\sqrt{(\Sigma^{1}_{Z})_{jj}}$ is sub-Gaussian with parameter $\sigma_{Z}^{1}$ for $j=1,\dots,p$.
@@
$||\beta^{*}||_{1}$ is bounded by $M>0$.
@@
$||\hat{\beta}-\beta^{*}||_{1}$ $\leq\textsf{Const}\text{ }s^{*}\lambda_{n},$
with probability at least $1-c_{1}\exp{\left(-c_{2}n\lambda_{n}^{2}\right)}$ for some constants $\textsf{Const},c_{1},c_{2}>0$. Further, if
$\min_{j\in S^{*}}|\beta^{*}_{j}|\gg\textsf{Const}\text{ }s^{*}\lambda_{n}\quad\text{as }n\xrightarrow[]{}\infty,$
and $\hat{S^{*}}$ is defined using the hard-threshold Lasso estimator as
$\hat{S}^{*}=\{j\text{ }|\text{ }|\hat{\beta}_{j}|>\textsf{Const}\text{ }s^{*}\lambda_{n}\},$
@@
Note that $s^{*}\lambda_{n}=o(1)$ implies $s^{*}\ll\sqrt{n}$. Consider the classical setting, where the number of samples $n$ goes to infinity, but the sparsity index $s^{*}$ do not depend on $n$. One could choose $\lambda_{n}^{2}=1/n^{1-\delta}$ for some $\delta>0$. Then with probability greater than $1-\tilde{c}_{1}\exp(-\tilde{c}_{2}n^{\delta})$, which converges to 1 as $n\xrightarrow[]{}\infty$, we have an error bound for $||\hat{\beta}-\beta^{*}||_{1}$ that converges to 0.

### v2

@@ 4.3 Incorporation of pre-treatment covariates
Lin’s estimator [Lin, 2013] improves the CUPED estimator by running separate regressions in the treated and control groups, with the agnostic property that it does not induce higher variance than the difference-in-means estimator without any model assumptions. Lin’s estimator estimates $\tau$ as
@@
To adapt the baseline approach to use pre-treatment covariates, instead of the simple difference-in-means estimator, one can use covariate-adjusted difference-in-means estimators. For either $\text{method}\in\{\text{CUPED},\text{Lin}\}$, we consider the following objective:
$\hat{\beta}=\arg\min_{\beta}\Bigg{|}\Bigg{|}\left(\frac{{\hat{\tau}_{\text{method}}}_{j}}{{\sqrt{\hat{\Sigma}_{\text{method}}}_{jj}}}\right)_{j=1}^{p}-\beta\text{ }\Bigg{|}\Bigg{|}_{2}^{2}\quad\text{subject to }||\beta||_{0}\leq s_{\tau}.$ (16)
@@
With these definitions, inference proceeds again as in Section 3 using the second split with covariate-adjusted $\tilde{Y}$, but restricted to the set $\hat{S}_{\tau}$.
It is possible to show under appropriate regularity and model assumptions that this baseline approach can consistently recover $S_{\tau}$. (For example, assumptions similar to those we make in our subsequent theoretical development in Section 5.2 would suffice.)
However, this approach may be quite sample inefficient in recovering the appropriate subset in the low signal-to-noise-ratio (SNR) regimes. This is because the baseline approach carries out detection independently in each dimension, without leveraging structure jointly across the dimensions. We show later via synthetic and semi-synthetic simulation results that the baseline approach exhibits a failure to select the right subset in the low SNR regime (see Figure 3 and 4 in Section 6). Notably, the examples discussed in Section 1 generally fall in these low SNR regimes. Motivated by this challenge, in the next section we take a sparse-linear-regression-based approach to subset selection, using the Lasso, which is known to generally perform well in low SNR regimes, cf. Hastie et al. [2020].
@@ 5 Our Proposed Approach:
Our approach starts from the following observation: Rather than considering each individual outcome dimension separately, they can be considered as a group as represented by the Hotelling’s $T^{2}$ test statistic in (6). This leads us to consider the following objective:
@@
where $\hat{\tau}^{S}_{\text{DiM}}$, $\hat{\Sigma}_{\text{DiM}}^{S}$ on the right-hand side are computed for each set $S\subset\{1,\dots,p\}$.
In the next subsection, we show that, in fact, the preceding problem is equivalent to selecting $\hat{S}_{\tau}=\{j|\hat{\beta}_{j}\neq 0\}$ where $\hat{\beta}$ is estimated by fitting a weighted linear regression with carefully chosen weights $W_{i}$ as
$\hat{\beta}=\arg\min_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta})^{2}\quad\text{subject to }||\beta||_{0}\leq s_{\tau}.$ (18)
Again, $||\beta||_{0}=\sum_{i=1}^{p}1\{\beta_{i}\neq 0\}$ is the $\ell_{0}$ norm of $\beta$. Note that we have treatment assignments as response variables and centered outcomes as explanatory variables. Here, $\bar{Y}$ is the mean of outcomes on the pooled data.
The objective in (17) represents a combinatorially hard problem in $p$, which we assume is large. Of note, the objective in (18) still presents computational complexity challenges as it does not yield a closed-form solution as did the baseline objective of (10). However, by reformulating (17) as a weighted linear regression with subset selection (18), we can employ various computationally efficient statistical tools that have been developed for the regression setting, such as forward subset selection, backward subset selection, or the Lasso. In the following, we focus on Lasso, which solves a convex relaxation of (18) by replacing the $\ell_{0}$ norm with the $\ell_{1}$ norm:
$\hat{\beta}=\arg\min_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta})^{2}\quad\text{subject to }||\beta||_{1}\leq t,$
where $||\beta||_{1}=\sum_{i=1}^{p}|\beta_{i}|$ is the $\ell_{1}$ norm of $\beta$, and $t\geq 0$ is a tuning parameter. Due to convex duality, the above problem is equivalent to the more common penalized regression form
$\hat{\beta}=\arg\min_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta})^{2}+2\lambda||\beta||_{1},$ (19)
@@
Having identified a subset $\hat{S}_{\tau}$ on the first data split, inference proceeds exactly as in Section 4.2.
@@
$(\hat{\beta},\hat{\alpha})=\arg\min_{\beta,\alpha}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta}-(X_{i}-\overline{X})^{\intercal}{\alpha})^{2}+2\lambda||\beta||_{1}.$ (20)
Note that
$\arg\min_{\beta}\min_{\alpha}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}{\beta}-(X_{i}-\overline{X})^{\intercal}{\alpha})^{2}$
$=\arg\min_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(\tilde{Y}-\bar{\tilde{Y}})^{\intercal}\beta)^{2},$
@@
Having identified $\hat{S}_{\tau}$, using the second split, we compute the covariate-adjusted difference-in-means estimator such as the CUPED estimator or Lin’s estimator on $\hat{S}_{\tau}$. The inference proceeds again as in Section 3 with covariate-adjusted $\tilde{Y}$ restricted to the set $\hat{S}_{\tau}$.
Our method shares a close relationship with the projection test [Lopes et al., 2011, Huang, 2015, Liu et al., 2022] in high-dimensional mean testing, as both involve finding sparse $\beta$ using a quadratic objective function with $\ell_{1}$ penalization. We focus on causal inference, and in particular as a result, we also consider pre-treatment covariates. Moreover, we employ a linear regression specification, making it easier for practitioners to implement our method with tools like glmnet in R.
Algorithm 1 provides a summary of the proposed method given two sets of data. As mentioned before, our data may naturally be divided in clinical settings by a POC study and a larger clinical trial after. Alternatively, we can employ a sample-splitting approach with a single pool of data. For a sample-splitting approach, the results are often sensitive to the arbitrary choice of a one-time random split of the data [Meinshausen et al., 2008]. Therefore, it is recommended to employ multiple splits and aggregate p-values from each split. In Appendix C, we present how we apply multi-sample splitting results from Meinshausen et al. [2008] to our proposed method when two data sets are obtained through sample-splitting.
Input: $\mathcal{D}_{1}=\{(T_{i},X_{i},Y_{i})\}_{i=1}^{n_{1}}$ and $\mathcal{D}_{2}=\{(T_{i},X_{i},Y_{i})\}_{i=1}^{n_{2}}$
@@ 1 Using the first split $\mathcal{D}_{1}$, estimate
where $\tilde{Y}_{ij}$ is defined as in (14) or (15). Its variance estimate $\hat{\Sigma}$ can be obtained using (4) with $\tilde{Y}$ restricted to the set $\hat{S}_{\tau}$.
Algorithm 1 Subset Selection via Sparse Regression (Single-Split)
In the remainder of the section, we first discuss the reduction of (17) to a weighted least squares regression; we then present a result that provides conditions under which we recover $S_{\tau}$.
@@ 5.1 Reduction to Weighted Linear Regression
In the following, we show that the weighted least-squares problem in (18) is equivalent to maximization of the Hotelling $T^{2}$ statistic in (17), which motivates the use of sparse linear regression in our approach.
Consider the following weighted least squares regression:
$\hat{\beta}=\arg\min_{\beta}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}^{S}-\overline{Y^{S}})^{\intercal}{\beta})^{2},$ (22)
where $Y_{i}^{S}$ is an outcome vector restricted to the given set $S$ as $Y_{i}^{S}:=(Y_{i,j}|j\in S)^{\intercal}.$ We set the weights $W_{i}$ so that the least-squares problem becomes equivalent to a signal strength maximization problem, where the signal strength is defined using Hotelling $T^{2}$ statistic in (6).
@@
On the right-hand side, ${\hat{\tau}_{\text{DiM}}^{S}}$ and $\hat{\Sigma}_{\text{DiM}}^{S}$ are obtained for each $S$. Note that the right-hand side is the Hotelling $T^{2}$ statistic in (6). This means that with our chosen weights, the weighted least-squares problem in (18) becomes equivalent to maximization of the Hotelling $T^{2}$ statistic as in (17), as desired.
@@ 5.2 Recovery Results for the Lasso
Recall that $\hat{\beta}$ is defined as the solution of the empirical risk:
$(\hat{\beta},\hat{\alpha})=\arg\min_{\beta\in\mathbb{R}^{p},\alpha\in\mathbb{R}^{m}}\frac{1}{n}\sum_{i}W_{i}(T_{i}-(Y_{i}-\overline{Y})^{\intercal}\beta-(X_{i}-\overline{X})^{\intercal}\alpha)^{2}+2\lambda_{n}||\beta||_{1},$ (29)
@@
Define the optimal solution of the population risk, $\beta^{*}$, as
$\beta^{*}=\arg\min_{\beta\in\mathbb{R}^{p}}\min_{\alpha\in\mathbb{R}^{m}}\mathbb{E}\Bigg{[}\frac{1}{n}\sum_{i}W_{i}^{*}(T_{i}-(Y_{i}-\mathbb{E}[Y])^{\intercal}\beta-(X_{i}-\mathbb{E}[X])^{\intercal}\alpha)^{2}\Bigg{]},$ (30)
where $W_{i}^{*}=\frac{1}{\pi^{2}}T_{i}+\frac{1}{(1-\pi)^{2}}(1-T_{i})$. Then, we have
$\beta^{*}=\arg\min_{\beta\in\mathbb{R}^{p}}\mathbb{E}\Bigg{[}\frac{1}{n}\sum_{i}W_{i}^{*}(T_{i}-(Z_{i}-\mathbb{E}[Z])^{\intercal}\beta)^{2}\Bigg{]},$
where $Z_{i}$ is the (population-level) covariate-adjusted outcome defined as
$Z_{i}:=Y_{i}-(1-\pi)\cdot\left[\text{Var}(X)^{-1}\text{Cov}(X,Y(1))\right]^{\intercal}X_{i}-\pi\cdot\left[\text{Var}(X)^{-1}\text{Cov}(X,Y(0))\right]^{\intercal}X_{i}.$
Note that this adjustment of outcomes $Y$ on pre-treatment covariates $X$ is very similar to pre-treatment covariates adjustments with linear regressions in Lin’s estimator [Lin, 2013]. If pre-treatment covariates are not available, we simply set $Z=Y$.
The closed form of $\beta^{*}$ can be verified to be:
@@
By the Sherman-Morrison formula, we have
@@
In the following, we provide an error bound for $||\hat{\beta}-\beta^{*}||$, and show that $\hat{S}_{\tau}=\{j|\hat{\beta}_{j}\neq 0\}$ recovers the set $S^{*}:=\{j|\beta^{*}_{j}\neq 0\}$ asymptotically. However, note that even if the treatment effect is sparse—i.e., even if $|S_{\tau}|<p$ where $S_{\tau}=\{j|\tau_{j}\neq 0\}$ in (9)—it does not follow that $\beta^{*}$ is sparse. To ensure $\beta^{*}$ is sparse, we additionally require the inverse of $\Sigma_{Z}$ to be sparse. Below, we provide examples of structural causal models [Bollen, 1989, Pearl, 2009] where the inverse of $\Sigma_{Z}$ is sparse, resulting in a sparse $\beta^{*}$, with the set $S^{*}$ containing the target set $S_{\tau}$. We formalize this as the following sparsity assumption on $\beta^{*}$ for future reference. We also provide simulation results when the sparsity assumption is slightly violated in the Appendix D.
@@
Assume that $\epsilon_{1},\dots,\epsilon_{p}$ are mean zero and uncorrelated with each other. Then, we have $S_{\tau}=S^{*}$.
@@
In this example we consider the same linear structural causal model as in the previous example, but instead of assuming that $\epsilon_{1},\dots,\epsilon_{p}$ are uncorrelated with each other, we instead assume a cluster structure: there exist clusters $C_{1},\dots,C_{c}$ such that if $j\in C_{i}$ and $j^{\prime}\in C_{i^{\prime}}$ with $i\neq i^{\prime}$, then $\epsilon_{j}$ and $\epsilon_{j^{\prime}}$ are uncorrelated. Then, when $S_{\tau}\subseteq C_{1}$, we have
$S_{\tau}\subseteq S^{*}\subseteq C_{1}.$
With a sparse cluster structure, $s_{\tau}\leq s^{*}\leq|C_{1}|\ll p$.
@@
Continue to assume the same linear structural causal model, but now assume that $\epsilon$ follows a multivariate normal distribution and “most” variables are conditionally independent of each other given the remaining variables. For example, we could consider an autoregressive model (AR($k$)) for time-series data. Let
$S_{\tau}\subseteq C:=\{j\text{ }|\text{ }\exists i\in S_{\tau}\text{ such that}\epsilon_{i}\text{ and }\epsilon_{j}\text{ are not conditionally independent.}\}$. }
Then, the inverse covariance matrix $\Sigma_{Z}$ is sparse. Specifically, if $i\in S_{\tau}$ and $j\notin C$, $(\Sigma_{Z}^{-1})_{i,j}=0$. Therefore, we have
$S_{\tau}\subseteq S^{*}=C.$
With a sparsely connected graph structure for $\epsilon$, $s^{*}\ll p$.
For our theoretical results, we assume sub-Gaussian designs for $X$, $Y(0)$, $Y(1)$, and thus for $Z(0),Z(1)$ as well. Then, we make additional assumptions about their moments and the invertibility of the covariance matrix of $X$.
@@
$X_{1},\dots,X_{n}$ are mutually independent and identically distributed sub-Gaussian vectors with mean $\mu_{X}$ and covariance $\Sigma_{X}$. Moreover, $Y_{1}(0),\dots,Y_{n}(0)$ and $Y_{1}(1),\dots,Y_{n}(1)$ are mutually independent and identically distributed sub-Gaussian vectors with mean $\mu^{0}_{Y}$ and $\mu^{1}_{Y}$ respectively and covariance $\Sigma^{0}_{Y}$ and $\Sigma^{1}_{Y}$ respectively. Additionally, $Z_{1}(0),\dots,Z_{n}(0)$ and $Z_{1}(1),\dots,Z_{n}(1)$ are also mutually independent and identically distributed sub-Gaussian vectors. Let the mean be $\mu^{0}_{Z}$ and $\mu^{1}_{Z}$ respectively and the covariance be $\Sigma^{0}_{Z}$ and $\Sigma^{1}_{Z}$ respectively. More specifically, each variable $(Z(0))_{j}/\sqrt{(\Sigma^{0}_{Z})_{jj}}$ is sub-Gaussian with parameter $\sigma_{Z}^{0}$ and $(Z(1))_{j}/\sqrt{(\Sigma^{1}_{Z})_{jj}}$ is sub-Gaussian with parameter $\sigma_{Z}^{1}$ for $j=1,\dots,p$.
@@
There exists $M>0$ such that $||\beta^{*}||_{1}\leq M$.
@@
$||\hat{\beta}-\beta^{*}||_{1}$ $\leq c_{3}\text{ }s^{*}\lambda_{n},$
with probability at least $1-c_{1}\exp{\left(-c_{2}n\lambda_{n}^{2}\right)}$ for some constants $c_{1},c_{2},c_{3}>0$. Further, if
$\min_{j\in S^{*}}|\beta^{*}_{j}|\gg c_{4}s^{*}\lambda_{n}\quad\text{as }n\xrightarrow[]{}\infty,$
for some constant $c_{4}>0$ and $\hat{S^{*}}$ is defined using the hard-threshold Lasso estimator as
$\hat{S}^{*}=\{j\text{ }|\text{ }|\hat{\beta}_{j}|>c_{4}s^{*}\lambda_{n}\},$
@@
Consider the classical setting, where the number of samples $n$ goes to infinity, but the sparsity index $s^{*}$ do not depend on $n$. One could choose $\lambda_{n}^{2}=1/n^{1-\delta}$ for some $\delta>0$. Then with probability greater than $1-\tilde{c}_{1}\exp(-\tilde{c}_{2}n^{\delta})$, which converges to 1 as $n\xrightarrow[]{}\infty$, we have an error bound for $||\hat{\beta}-\beta^{*}||_{1}$ that converges to 0.